# Gravitational lensing by charged black hole with global monopole in strong field limit

### v1

@@ IV Discussion
In this paper, we have investigated the gravitational lensing near a charged black hole with a global monopole in the strong field limit, focusing on the interplay between charge and the global monopole on lensing observables. Notably, the phenomena described here, such as the non-monotonic behavior of the angular separation, are particularly intriguing when both charge and the global monopole’s deficit angle are large, with critical values identified for both parameters (refer to the critical values for $\eta$ and $\epsilon$ in the last section). However, astrophysical black holes are expected to be nearly neutral, as any significant charge would likely be neutralized by surrounding matter. Furthermore, the deficit angle caused by a global monopole is expected to be extremely small. In typical grand unification scenarios, the parameter $\gamma$ in the metric (3.1) characterizing the global monopole is on the order of $\sim 10^{16}$ GeV Barriola and Vilenkin (1989), leading to a deficit angle $\Delta=32\pi^{2}\gamma^{2}\sim 4\pi\times 10^{-5}$ rad, which is negligibly small. Consequently, the phenomena described in this paper are unlikely to be observed in typical astrophysical black holes.
@@
Nevertheless, there is growing interest in analogue gravity systems Barcelo et al. (2005) as platforms for studying gravitational effects. These systems, which use optical metamaterials or other artificial materials to mimic the behavior of spacetime, allow for more flexibility in tuning parameters such as charge and deficit angle, which may be too small in natural systems. In particular, the theory of transformation optics Leonhardt (2006); Pendry et al. (2006) establishes a correspondence between the metric tensor and the permittivity and permeability tensors of materials, enabling the design of optical metamaterials to mimic astronomical phenomena. For instance, gravitational lensing effects have been experimentally visualized in such systems Sheng et al. (2013). In such optical metamaterials, the effective charge and deficit angle are not constrained to be small. In particular, by using an artificial waveguide bounded with rotational metasurface, the nontrivial effects of a spacetime topological defect have been experimentally emulated, and photon deflection in the topological waveguide has been observed Sheng et al. (2018). In this context, the phenomena predicted here may potentially be observed in analogue systems.
@@ V Conclusion
In this paper, we have investigated the phenomenon of gravitational lensing near a charged black hole with a global monopole, with particular focus on the effects of the black hole charge and the global monopole on key lensing observables. We have derived the coefficients $\bar{a}$ and $\bar{b}$ that characterize the deflection angle of light. Based on these results, we have investigated the effects of the global monopole and the black hole charge on two key observables in gravitational lensing: the angular separation $s$ between the first image and subsequent images, and the flux ratio $\mathcal{R}$, defined as the ratio of the flux intensity from the first image to the cumulative flux intensity of all other images. We have derived analytical expressions for the observables $s$ and $\mathcal{R}$, and performed both analytical and numerical analyses. The findings are summarized as follows:
For the angular separation $s$, we find that regardless of the value of the charge, $s$ always decreases as the deficit angle caused by the global monopole decreases. However, the behavior of $s$ as a function of the charge is highly dependent on the global monopole parameter. We identified three critical values of the global monopole parameter, which divide the parameter space into four distinct regions. In these regions, the angular separation $s$ can either increase or decrease monotonically with charge, or exhibit more complex behaviors such as an initial decrease followed by an increase, or a non-monotonic behavior with both a maximum and a minimum. These behaviors contrast sharply with the monotonic increase in angular separation as a function of charge observed in the absence of a global monopole.
@@
These results demonstrate that in the strong field limit, the charge and global monopole interact in complex ways, and their combined effects on the lensing observables cannot be attributed to the independent contributions of each factor. The resulting lensing signatures provide a rich landscape for further exploration.
While the charge and global monopole deficit angle are expected to be small in typical astrophysical black holes, these parameters are not constrained in analogue gravity systems. Therefore, the phenomena predicted here, including the non-monotonic behavior of the angular separation and flux ratio, may potentially be observed in optical metamaterials or other experimental setups designed to mimic black hole physics. These findings open up new avenues for testing gravitational lensing effects in laboratory-based systems, offering the possibility of future experimental verification of our theoretical predictions.
@@ Acknowledgments
Appendix A The derivative of angular separation with respect to $\eta$
$s(1,\eta)$ is given by:
@@ Appendix A The derivative of angular separation with respect to $\eta$
The numerical solution of $F(\epsilon_{1})=0$ gives $\epsilon_{1}\approx 0.182669$. To illustrate the behavior of $F(\epsilon)$ around $\epsilon_{1}$, we present a plot of $F(\epsilon)$ in Fig. 7. From Fig. 7, it is evident that $F(\epsilon)>0$ when $\epsilon<\epsilon_{1}$, while $F(\epsilon)<0$ when $\epsilon>\epsilon_{1}$.

### v3

@@ IV Discussion
In this paper, we have investigated the gravitational lensing near a charged black hole with a global monopole in the strong field limit, focusing on the interplay between charge and the global monopole on lensing observables. Notably, the phenomena described here, such as the nonmonotonic behavior of the angular separation, are particularly intriguing when both charge and the global monopole’s deficit angle are large, with critical values identified for both parameters (refer to the critical values for $\eta$ and $\epsilon$ in the last section). It is interesting to examine whether the effects of charge and the global monopole on lensing observables can be detected for reasonable values of the global monopole parameter and black hole charge in astrophysical settings. To this end, we consider black holes at the centers of three galaxies, i.e., the Milky Way, M87, and M104, as gravitational lenses. The relevant parameters are $M=4.3\times 10^{6}M_{\odot}$ and $D_{\text{OL}}=8.3\text{kpc}$ for Sgr A* Abuter et al. (2023), $M=6.5\times 10^{9}M_{\odot}$ and $D_{\text{OL}}=16.8\text{Mpc}$ for M87* Akiyama et al. (2019), and $M=1.0\times 10^{9}M_{\odot}$ and $D_{\text{OL}}=9.55\text{Mpc}$ for M104* Kormendy et al. (1996); McQuinn et al. (2016). Additionally, we assume that the lens is located between the light source and the observer, such that $D_{\text{LS}}=D_{\text{OL}}=D_{\text{OS}}/2$. The results for the angular separation $s$, the flux ratio $\mathcal{R}$, the magnification of the first image $\mu_{1}$, the angular radius of the first Einstein ring $\theta_{1}^{E}$, and the time delay between the first and second images $\Delta T_{2,1}^{s+}$ are presented in Tables. 1-5. These numerical results indicate that, for astrophysical black holes, the influence of charge and global monopoles on lensing observables is negligible. Even with exaggerated charge and deficit angle due to the global monopole ($\eta=0.01$ and $\Delta=4\pi\times 10^{-3}$, or equivalently $\epsilon=1-\frac{\Delta}{4\pi}=0.999$), the results remain nearly indistinguishable from those of Schwarzschild black holes.
@@
In fact, astrophysical black holes are expected to be nearly neutral, as any significant charge would likely be neutralized by surrounding matter. Furthermore, the deficit angle caused by a global monopole is expected to be extremely small. In typical grand unification scenarios, the parameter $\gamma$ in the metric (2.1) characterizing the global monopole is on the order of $\sim 10^{16}$ GeV Barriola and Vilenkin (1989), leading to a deficit angle $\Delta=32\pi^{2}\gamma^{2}\sim 4\pi\times 10^{-5}$ rad, which is negligibly small. Consequently, the rich gravitational lensing phenomena described in this paper are unlikely to be observed in typical astrophysical black holes.
Schwarzschild Schwarzschild with global monopole RN RN with global monopole
Sgr A* 0.0333 0.0335 0.0335 0.0337
M87* 0.0248 0.0250 0.0250 0.0251
M104* 0.0067 0.0068 0.0068 0.0068
Table 1: The angular separation $s$ (in microarcsecond) between the first image and the other images for different black holes as lenses. The charge and global monopole parameters are taken as $\eta=0.01$ and $\epsilon=0.999$, respectively. The same parameters apply to all following tables.
Schwarzschild Schwarzschild with global monopole RN RN with global monopole
Sgr A* 535.16 533.48 531.42 529.75
M87* 535.16 533.48 531.42 529.75
M104* 535.16 533.48 531.42 529.75
Table 2: The flux ratio $\mathcal{R}$ for different black holes as lenses.
Schwarzschild Schwarzschild with global monopole RN RN with global monopole
Sgr A* 4.7688 4.8034 4.7843 4.8189
M87* 2.6597 2.6790 2.6684 2.6877
M104* 0.1948 0.1962 0.1954 0.1969
Table 3: The magnification of the first image $\mu_{1}\times 10^{22}$ for different black holes as lenses.
Schwarzschild Schwarzschild with global monopole RN RN with global monopole
Sgr A* 26.613 26.653 26.569 26.609
M87* 19.875 19.905 19.842 19.872
M104* 5.379 5.387 5.370 5.378
Table 4: The angular radius of the first Einstein ring $\theta_{1}^{E}$ (in microarcsecond) for different black holes as lenses.
Schwarzschild Schwarzschild with global monopole RN RN with global monopole
Sgr A* 0.1953 0.1956 0.1950 0.1953
M87* 295.2550 295.7090 294.7790 295.2330
M104* 45.4239 45.4938 45.3507 45.4205
Table 5: The time delay between the first and second images $\Delta T_{2,1}^{s+}$ (in hour) for different black holes as lenses.
@@
Nevertheless, there is growing interest in analog gravity systems Barcelo et al. (2005) as platforms for studying gravitational effects, particularly gravitational lensing Fischer and Visser (2002); Sheng et al. (2013). These systems, which use acoustic, optical, or other artificial materials to mimic the behavior of spacetime, allow for more flexibility in tuning parameters such as charge and deficit angle, which may be too small in natural systems. In particular, the theory of transformation optics Leonhardt (2006); Pendry et al. (2006) establishes a correspondence between the metric tensor and the permittivity and permeability tensors of materials, enabling the design of optical metamaterials to mimic astronomical phenomena. For instance, gravitational lensing effects have been experimentally visualized in such systems Sheng et al. (2013). In such optical metamaterials, the effective charge and deficit angle are not constrained to be small. In particular, by using an artificial waveguide bounded with rotational metasurface, the nontrivial effects of a spacetime topological defect have been experimentally emulated, and photon deflection in the topological waveguide has been observed Sheng et al. (2018). In this context, the phenomena predicted here may potentially be observed in analog systems.
@@ V Conclusion
In this paper, we have investigated the phenomenon of gravitational lensing near a charged black hole with a global monopole, with particular focus on the effects of the black hole charge and the global monopole on key lensing observables. We have derived the coefficients $\bar{a}$ and $\bar{b}$ that characterize the deflection angle of light. Based on these results, we have investigated the effects of the global monopole and the black hole charge on key observables in gravitational lensing, including the angular separation, the flux ratio, the magnification, the angular radius of the Einstein ring, and the time delay. The findings are summarized as follows:
For the angular separation $s$, we find that regardless of the value of the charge, $s$ always decreases as the deficit angle caused by the global monopole decreases. However, the behavior of $s$ as a function of the charge is highly dependent on the global monopole parameter. We identified three critical values of the global monopole parameter, which divide the parameter space into four distinct regions. In these regions, the angular separation $s$ can either increase or decrease monotonically with charge, or exhibit more complex behaviors such as an initial decrease followed by an increase, or a nonmonotonic behavior with both a maximum and a minimum. These behaviors contrast sharply with the monotonic increase in angular separation as a function of charge observed in the absence of a global monopole.
@@
Similarly, for the magnification of the first relativistic image $\mu_{1}$, we find that irrespective of the charge, $\mu_{1}$ consistently decreases as the global monopole parameter increases. However, the dependence of $\mu_{1}$ on the charge varies with the global monopole parameter. Specifically, two critical values of the global monopole parameter determine whether $\mu_{1}$ increases or decreases monotonically with charge, or whether it initially decreases and then increases.
These results demonstrate that in the strong field limit, the charge and global monopole interact in complex ways, and their combined effects on the lensing observables cannot be attributed to the independent contributions of each factor. The resulting lensing signatures provide a rich landscape for further exploration. In addition, we numerically analyzed the angular size of the first Einstein ring $\theta_{1}^{E}$, and the time delay between the first and second images $\Delta T_{2,1}^{s+}$ as functions of the charge and global monopole parameters. Our findings indicate that both observables decrease with increasing global monopole parameter $\epsilon$ and increasing charge.
While the charge and global monopole deficit angle are expected to be small in typical astrophysical black holes, these parameters are not constrained in analog gravity systems. Therefore, the phenomena predicted here, including the nonmonotonic behavior of the angular separation and flux ratio, may potentially be observed in optical metamaterials or other experimental setups designed to mimic black hole physics. These findings open up new avenues for testing gravitational lensing effects in laboratory-based systems, offering the possibility of future experimental verification of our theoretical predictions.
@@ Acknowledgments
Appendix A DERIVATIVE OF ANGULAR SEPARATION WITH RESPECT TO $\eta$
$s(1,\eta)$ is given by
@@ Appendix A DERIVATIVE OF ANGULAR SEPARATION WITH RESPECT TO $\eta$
The numerical solution of $F(\epsilon_{1})=0$ gives $\epsilon_{1}\approx 0.1826688$. To illustrate the behavior of $F(\epsilon)$ around $\epsilon_{1}$, we present a plot of $F(\epsilon)$ in Fig. 11. From Fig. 11, it is evident that $F(\epsilon)>0$ when $\epsilon<\epsilon_{1}$, while $F(\epsilon)<0$ when $\epsilon>\epsilon_{1}$.